OpenStudy (anonymous):

How would I multiply and simplify (if possible)? √7(√(x) - 7√7)

OpenStudy (abb0t):

\[\sqrt{7} = 7^{\frac{ 1 }{ 2 }} \]

OpenStudy (abb0t):

\[7^{\frac{ 1 }{ 2 }} (x^{\frac{ 1 }{ 2 }} - 7(7^{\frac{ 1 }{ 2} }))\] distribute and add the exponents.

OpenStudy (anonymous):

Thanks but could you explain a little more in how I could multiply this step by step?

OpenStudy (abb0t):

distribute the \[7^{\frac{ 1 }{ 2 }}\] across. Remember, when you multiply exponents, you are adding them. for example: \[x^a(x^{100a}) = x^{a+100a} = x^{101a}\]

OpenStudy (abb0t):

Another example: \[7^\frac{ 1 }{ 3 }(b^{\frac{ 1 }{ 3 }}) = \sqrt[3]{7b} = 7b^{\frac{ 1 }{ 3 }}\]

OpenStudy (anonymous):

Oh I think I get it now thnx. :D